OpenStudy (camzzzie):

I'll medal for help!!! I need to find a b and c but I always get messed up with the method :( 1.39= a(0)^2+b(0)+c 1.26= a(18)^2+b(18)+c 0= a(45)^2+b(45)+c

OpenStudy (anonymous):

\[ 1.39= a(0)^2+b(0)+c \]This simplifies to: \[ 1.39=c \]

OpenStudy (camzzzie):

then plug c in?

OpenStudy (anonymous):

Yes

OpenStudy (camzzzie):

ok . How do you find just a or b when they are both in the equation?

OpenStudy (anonymous):

Isolate one of the variables.

OpenStudy (camzzzie):

0= a(45)^2+b(45)+ 1.39 -1.39-b(45)=a (45)^2

OpenStudy (camzzzie):

is that right?

OpenStudy (anonymous):

To isolate, you need to move the other variable to the other side of the equation.

OpenStudy (anonymous):

If you were to isolate a, for example, you would need to move over b, and then divide by the coefficient (number next to it) of a.